Question 2 A single point charge sits alone in a region of space. The electric field due to the charge at a distance of 0.283 meters is 8.19e+3 N/C. Calculate the magnitude of the charge on the point charge.
Answer #1

given

electric field E = 8.19*10^3 N/C

distance r = 0.283 m

from the relation

electric field due to a point charge is E = kq/r^2

here k = 9*10^9

q = Er^2/k

q = 8.19*10^3*0.283^2/9*10^9

q = 7.28*10^-8 C